Find the exponential function f(x)=a^x whose graph goes through the point (3,1/125).
a=
I have no idea how to go about finding the equation :S
Answer #1
graph coordinates are (x , y) or (x , f(x)) 1/125 = a^3 5 is the cube root of 125, so ___ 1/5 = a
answered by: Aryna